OpenStudy (anonymous):

S = 2*pi*d*sqrt(d^2+h^2) rearrange to make h the subject

OpenStudy (anonymous):

Square everything first...

OpenStudy (anonymous):

would I not divide everything by 2*pi*d first?

OpenStudy (anonymous):

Ok, do that first...

OpenStudy (anonymous):

S/(2*pi*d)=sqrt(d^2+h^2)

OpenStudy (anonymous):

Square everything...

OpenStudy (anonymous):

(S/(2*pi*d))^2=d^2+h^2

OpenStudy (anonymous):

Subtract d^2 and then square root...

OpenStudy (anonymous):

(S/(2*pi*d))^2-d^2=h^2 h=sqrt((S/(2*pi*d))^2-d^2)

OpenStudy (anonymous):

Good enough?

OpenStudy (anonymous):

So is this my answer:

OpenStudy (anonymous):

Thank you very much! :)

OpenStudy (anonymous):

yw:-)